OpenStudy (awesome781):

Solve the following rational exponent equation m^3/4=27

OpenStudy (anonymous):

i'm not sure if it is 81

OpenStudy (anonymous):

Are there any multiple choice?

OpenStudy (anonymous):

Because I also got 81

OpenStudy (awesome781):

Nope but thnx

OpenStudy (shaik0124):

i will help

OpenStudy (awesome781):

Really......thnx

OpenStudy (shaik0124):

m3=27*4 m3=108 now solve

OpenStudy (shaik0124):

m^3*1/3=(108)^1/3 m=(108)^1/3

OpenStudy (shaik0124):

@Awesome781

OpenStudy (shaik0124):

m=36

OpenStudy (awesome781):

So 81 is wrong?

OpenStudy (shaik0124):

yup

OpenStudy (shaik0124):

how 81?

OpenStudy (shaik0124):

??

OpenStudy (awesome781):

@shaik0124 ......so the answer is not 36?

OpenStudy (anonymous):

I would help but this site dont let me draw.

OpenStudy (shaik0124):

y i am also unable to draw y like that@timo

OpenStudy (shaik0124):

(108)^0.33=108*0.33=891/25=35.64 which is almost equal to 36 so 36 is answer

OpenStudy (awesome781):

@timo86m did u get the same answer as @shaik0124 ?

OpenStudy (anonymous):

i did not even try it. OS wont let me use the draw button.

OpenStudy (awesome781):

That's ok

OpenStudy (anonymous):

should be 81 tho